Consider the upward motion of a particle under gravity with a velocity of projection uo and resistance mkv2. Show that the velocity V at the time t and distance x from point lf projection are related as

2gx/Vt2 = ln[(uo2+Vt2)/(V2+Vt2)], where k=g/Vt2

1
Expert's answer
2021-06-07T15:33:34-0400

Using Newton’s 2nd law gives:


"m\dfrac{d^2x}{dt^2}=-mk(\dfrac{dx}{dt})^2-mg"

"\dfrac{dV}{dt}=-kV^2-g"

"\dfrac{dV}{dt}=\dfrac{dV}{dx}\cdot\dfrac{dx}{dt}=V\dfrac{dV}{dx}"

"V\dfrac{dV}{dx}=-kV^2-g"

"\dfrac{VdV}{kV^2+g}=-dx"

"\int \dfrac{VdV}{kV^2+g}=-\int dx"

"\int \dfrac{VdV}{kV^2+g}"

"z=kV^2+g, dz=2kVdV"


"VdV=\dfrac{1}{2}dz"


"\int \dfrac{VdV}{kV^2+g}=\int \dfrac{dz}{2kz}=\dfrac{1}{2k}\ln|z|+C_1"


"=\dfrac{1}{2k}\ln(kV^2+g)+C_1"


"\dfrac{1}{2k}\ln(kV^2+g)=-x+\dfrac{1}{2}C_2"

"2kx=-\ln(kV^2+g)+kC_2"

"t=0: 0=-\ln(ku_0^2+g)+kC_2"

"=>kC_2=\ln(ku_0^2+g)"

"2kx=\ln(\dfrac{ku_0^2+g}{kV^2+g})"

If "k=\dfrac{g}{V_t^2}"


"\dfrac{2gx}{V_t^2}=\ln(\dfrac{u_0^2+V_t^2}{V^2+V_t^2})"